OpenStudy (anonymous):

solve for x: 2/5(x-4)=2x

OpenStudy (campbell_st):

is the question \[\frac{2(x -4)}{5} = 2x\]

OpenStudy (anonymous):

no 2/5 as a fraction

OpenStudy (anonymous):

@Marethyu I got -1=x

OpenStudy (campbell_st):

is ( x - 4) in the denominator..?

OpenStudy (anonymous):

no^

OpenStudy (anonymous):

its 2over5(x-4)=2x

OpenStudy (anonymous):